# Imitative Membership Inference Attack

### v1

@@ 3.2 Discussion of $\mathsf{IMIA}$
Algorithm 3 Imitative Membership Inference Attack ($\mathsf{IMIA}$) in the adaptive setting. The adversary trains NN imitative in and imitative out models for the query instance to perform the attack.
@@
Non-parametric Modeling. State-of-the-art MIAs [3, 9] rely on parametric modeling to compute final scores. For instance, LiRA [3] estimates the likelihood ratio by fitting Gaussian distributions to scaled confidence scores. However, accurately estimating the parameters of these distributions requires training a large number of shadow models, and the performance degrades significantly when computational resources are constrained [9, 60]. To address this, $\mathsf{IMIA}$ adopts a non-parametric view and computes a score by measuring the squared distance between attack signals derived from imitative in and imitative out models. This design avoids the cost of parameter estimation, enabling $\mathsf{IMIA}$ to outperform state-of-the-art MIAs while using less than 5% of their computational cost.
Connections with Model Distillation. The imitative training procedure in $\mathsf{IMIA}$ is conceptually related to model distillation, a technique used in several prior MIAs [58, 15, 30, 34]. However, $\mathsf{IMIA}$ differs from these works in several key aspects: (i) Tailored loss function. While prior attacks employ standard KL divergence for distillation, $\mathsf{IMIA}$ introduces the imitation loss that is specifically designed to distill the structural logit information most indicative of membership. (ii) Dual behavior modeling. Previous MIAs that utilize model distillation only model the target’s behavior on non-members. In contrast, our two-phase imitative training first creates an imitative out model and then fine-tunes it on pivot data to create an imitative in model. This second phase explicitly encodes membership signals while preserving the target model’s knowledge, allowing the adversary to take advantage of both in and out behaviors to mount a more powerful attack. (iii) Significant performance gain. In our experiments, we observe that prior distillation-based MIAs achieve subpar performance. In contrast, $\mathsf{IMIA}$ significantly outperforms all existing attacks while requiring substantially less computational overhead.
@@ 4 Evaluation
We conduct a comprehensive evaluation of $\mathsf{IMIA}$ across multiple datasets and attack settings. We begin with the experimental setup in Section˜4.1, then evaluate performance, efficiency, and robustness in both settings in Section˜4.2. In Section˜4.3, we provide a quantitative assessment of the effectiveness of imitative training. We further analyze the impact of attack components in Section˜4.4 and present additional analyses and assess its performance against defenses in Section˜4.5.
@@ 4.1 Experimental Setup
Datasets. We use four benchmark image datasets (i.e., MNIST [28], Fashion-MNIST [55], CIFAR-10 [26], and CIFAR-100 [26]) for our main experiments. To demonstrate the generalizability of $\mathsf{IMIA}$, we also report attack results on two widely used non-image datasets (i.e., Purchase and Texas [48]) in Section˜4.5. A detailed description of all datasets is provided in Section˜A.1.
Network Architecture. We consider four widely used neural network architectures for image classification: ResNet50 [18], VGG16 [49], DenseNet121 [21], and MobileNetV2 [47]. We follow the training configurations used in prior work [3] to mitigate overfitting. Specifically, we use the SGD algorithm with a learning rate of $0.1$, momentum of $0.9$, and weight decay [27] set to $5\times 10^{-4}$. We also adopt a cosine learning rate schedule [38] for optimization and apply data augmentation techniques [61] during training. The training and validation accuracy of the target models are reported in Section˜A.2.
@@
  • LOSS [59] uses the loss of the query instance for inference.
  • Entropy [50] leverages a modified prediction entropy estimation for membership inference.
  • Calibration [53] employs a technique called difficulty calibration, which adjusts the query instance’s loss by calibrating it on shadow models.
  • Attack-R [58] compares the loss of the query instance on the target model to its losses on shadow models. The final score is the proportion of shadow models for which the target model yields a lower loss.
  • Attack-D [58] uses model distillation [20] to train shadow models and computes the same ratio scores as Attack-R.
  • SeqMIA [30] leverages model distillation to replicate the learning process of the target model, extracting various membership metrics from its loss trajectory for attack.
  • LiRA [3] adapts different strategies for two attack settings. In the adaptive setting, it trains both shadow in and shadow out models and applies a likelihood ratio test for inference. In the non-adaptive setting, it uses only shadow out models to conduct a one-sided hypothesis test.
  • Canary [54] improves upon LiRA by employing adversarial optimization to enhance inference performance.
  • GLiRA [15] enhances LiRA with model distillation to obtain shadow out models for a one-sided hypothesis test.
  • RMIA [60] performs multiple pairwise likelihood ratio tests between the query instance and randomly selected population instances to compute the score.
  • RAPID [19] trains a neural network for membership inference by combining raw and calibrated loss values [53].
  • PMIA [9] approximates the posterior odds test by using behaviors of the shadow model’s training set as proxies.
@@
Evaluation Procedures. We divide each dataset into two disjoint subsets: the query set $D_{\text{query}}$ and the auxiliary set $D_{\text{auxiliary}}$. The target model is trained on a randomly sampled subset of $D_{\text{query}}$. In the adaptive setting, we follow [3, 54] and provide the adversary with the same data as the query set to prepare their attack (i.e., $D^{\text{adapt}}_{\text{adv}}=D_{\text{query}}$). In the non-adaptive setting, consistent with [19, 9, 60], the adversary’s dataset is set to the auxiliary set (i.e., $D^{\text{nonadapt}}_{\text{adv}}=D_{\text{auxiliary}}$), ensuring that the adversary cannot access queries when training shadow models. Details about the data split are provided in Table˜19.
Evaluation Metrics. Following previous studies [3, 60, 54, 9, 19], we focus on evaluating the attack performance on the low false positive rate regime. Specifically, we use the following evaluation metrics for our experiments:
  • TPR@0.001%FPR. It directly reflects the extent of privacy leakage of the model by allowing only one (or a few) false positives to compute the true positive rate.
@@
Attack Setup. For baselines, we train the recommended number of shadow models as specified in their original papers to ensure their best performance. For instance, LiRA trains 256 shadow models to estimate the parametric likelihood ratio. (A complete list of shadow model counts for baselines is provided in Section˜A.3) In contrast, $\mathsf{IMIA}$ uses a fixed set of $N=10$ imitative models to highlight its computational efficiency. We train the imitative out models for $T_{1}=100$ epochs across all datasets, matching the number of epochs used in shadow training [3]. For the imitative in models, we continue training for $T_{2}=20$ using pivot data. All hyperparameters are set according to the original implementations, and we report the average metrics over ten runs as the final results.
Table 1: Performance comparison of non-adaptive attacks on ResNet50 across four image datasets (i.e., MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100). $\mathsf{IMIA}$ significantly outperforms the strongest baseline (underlined, PMIA [9]) while requiring less than 5% of its computational cost (see Section˜4.2 for efficiency comparison). The %Imp. indicates the relative improvement of $\mathsf{IMIA}$ compared to the strongest baseline. The best result is in bold.
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Table 2: Performance comparison of adaptive attacks on ResNet50 across four image datasets. $\mathsf{IMIA}$ outperforms the strongest baseline (underlined) while requiring less than 5% of its computational cost. The best result is in bold.
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@ 4.2 Evaluation of $\mathsf{IMIA}$
Performance in the Non-Adaptive Setting. We first evaluate $\mathsf{IMIA}$ in the non-adaptive setting, and the results on ResNet-50 are shown in Table˜1. $\mathsf{IMIA}$ consistently outperforms all existing methods across the evaluated metrics. For instance, on MNIST, it achieves a true positive rate (TPR) of 1.01% at a false positive rate (FPR) of 0.001%, which is at least three times as high as the best-performing baseline (i.e., PMIA). The advantage is even more pronounced on Fashion-MNIST, with a 2.52% TPR at the same FPR, over ten times higher than baselines. These results represent a substantial advancement, as such datasets were previously considered hard to attack. Although the improvement in balanced accuracy is relatively modest, this is consistent with prior studies [3, 58, 60]. Moreover, the consensus in the MIA community is that attacks should be evaluated using TPR at low FPR, where $\mathsf{IMIA}$ shows a clear advantage. Similar trends are observed on other model architectures; results and Receiver Operating Characteristic (ROC) curves are provided in Section˜B.4.
Performance in the Adaptive Setting. We also evaluate $\mathsf{IMIA}$ with SOTA MIAs in the adaptive setting. The results for ResNet-50 are shown in Table˜2, while performance on other architectures can be found in Section˜B.5. $\mathsf{IMIA}$ consistently achieves the best attack performance across all datasets, surpassing strong attacks like LiRA and Canary. For instance, $\mathsf{IMIA}$ achieves nearly double the performance of LiRA on CIFAR-10 at a TPR of 0.001% FPR, with similar improvements observed across other datasets.
Comparison with Distillation-based MIAs. While the imitative training used in $\mathsf{IMIA}$ is conceptually related to model distillation, it significantly outperforms distillation-based attacks (i.e., Attack-D, SeqMIA, and GLiRA) in both adaptive and non-adaptive settings. For instance, Attack-D yields negligible TPR at low FPRs, and in the non-adaptive setting, $\mathsf{IMIA}$ outperforms SeqMIA and GLiRA by at least a factor of five. We attribute this performance gap to key limitations in existing distillation-based MIAs: (i) These attacks directly apply the standard model distillation process [20], failing to capture the salient membership signals that are important for MIA. (ii) They use distillation to create shadow out models, which can only mimic the target’s behavior on non-member instances. In contrast, $\mathsf{IMIA}$ explicitly caputer membership signals by training both imitative in and out models, leveraging their behavioral discrepancies for a more effective attack.
@@
Efficiency Analysis. We measure the overall runtime of $\mathsf{IMIA}$ in both adaptive and non-adaptive settings on a cluster with a single A100 GPU, comparing it to leading MIAs. We do not include simple baselines (e.g., LOSS attack [59]) for comparison, as their performance is far inferior to $\mathsf{IMIA}$. As shown in Table˜3, $\mathsf{IMIA}$ requires substantially less computation. For instance, on CIFAR-10, the SOTA baselines (i.e., PMIA and LiRA) require 6 days to perform their attack, whereas $\mathsf{IMIA}$ completes in less than 7 hours, a reduction of more than 95%. This efficiency is consistently observed across all datasets.
@@
Figure 3: Performance on MNIST under varying computational budgets. $\mathsf{IMIA}$$_{\text{Gaussian}}$ employs a Gaussian likelihood ratio [3] on imitative models to compute the final scores.
@@
Performance Across Computational Budgets. To evaluate the trade-off between efficiency and attack effectiveness, we benchmark $\mathsf{IMIA}$ against leading MIAs using an A100 GPU in both non-adaptive and adaptive settings. We also introduce $\mathsf{IMIA}$$_{\text{Gaussian}}$, a variant that leverages the same imitative training as $\mathsf{IMIA}$ but employs parametric modeling for inference. Specifically, it follows LiRA [3] and fits the scaled confidence scores of imitative models as Gaussian distributions, and uses the likelihood ratio as the membership score.
@@ 4.3 Effectiveness of Imitative Training
Does Imitative Training Better Mimic the Target Model? We conduct a set of experiments comparing its ability to mimic the target model’s behavior against that of shadow training. Specifically, we train 256 shadow models and imitative models in the adaptive setting. We then collect the scaled confidence score distributions for both members and non-members by querying the shadow and imitative models. Prior studies [3, 54, 9] have empirically validated that these score distributions are well-approximated by a Gaussian distribution. Following this approach, we model each distribution as Gaussian and estimate its mean $\mu$ and standard deviation $\sigma$. Using these parameters, we compute the normalized residual of the target model’s true score $s$ as $r=\frac{s-\mu}{\sigma}$. If the models’ distributions truly replicate the target model’s behavior, the residuals $r$ for both members and non-members should follow a standard normal distribution, i.e., $r\sim N(0,1)$.

### v2

@@ 3.2 Discussion of $\mathsf{IMIA}$
Non-parametric Modeling. State-of-the-art MIAs [sp22lira, ndss26cpmia] rely on parametric modeling to compute final scores. For instance, LiRA [sp22lira] estimates the likelihood ratio by fitting Gaussian distributions to scaled confidence scores. However, accurately estimating the parameters of these distributions requires training a large number of shadow models, and the performance degrades significantly when computational resources are constrained [ndss26cpmia, icml24rmia]. To address this, $\mathsf{IMIA}$ adopts a non-parametric view and computes a score by measuring the squared distance between attack signals derived from imitative in and imitative out models. This design avoids the cost of parameter estimation, enabling $\mathsf{IMIA}$ to outperform state-of-the-art MIAs while using less than 5% of their computational cost.
Connections with Model Distillation. The imitative training procedure in $\mathsf{IMIA}$ is conceptually related to model distillation, a technique used in several prior MIAs [ccs22enhanced, tifs25glira, ccs24seqmia, ccs22trajectory]. However, $\mathsf{IMIA}$ differs from these works in several key aspects: (i) Tailored loss function. While prior attacks employ standard KL divergence for distillation, $\mathsf{IMIA}$ introduces the imitation loss that is specifically designed to distill the structural logit information most indicative of membership. (ii) Dual behavior modeling. Previous MIAs that utilize model distillation only model the target’s behavior on non-members. In contrast, our two-phase imitative training first creates an imitative out model and then fine-tunes it on pivot data to create an imitative in model. This second phase explicitly encodes membership signals while preserving the target model’s knowledge, allowing the adversary to take advantage of both in and out behaviors to mount a more powerful attack. (iii) Significant performance gain. In our experiments, we observe that prior distillation-based MIAs achieve subpar performance. In contrast, $\mathsf{IMIA}$ significantly outperforms all existing attacks while requiring substantially less computational overhead.
Algorithm 3 Imitative Membership Inference Attack ($\mathsf{IMIA}$) in the adaptive setting. The adversary trains NN imitative in and imitative out models for the query instance.
@@ 4.1 Experimental Setup
Datasets. We use four image datasets (i.e., MNIST [mnist], Fashion-MNIST [fmnist], CIFAR-10 [cifar10], and CIFAR-100 [cifar10]) for our main experiments. To demonstrate the generalizability of $\mathsf{IMIA}$, we also report attack results on two widely used non-image datasets (i.e., Purchase and Texas [sp17miashokri]) in Section˜4.5. The dataset descriptions are provided in Appendix A.1.
Network Architecture. We consider four widely used neural network architectures for image classification: ResNet [cvpr16resnet], VGG16 [vgg], DenseNet121 [densenet], and MobileNetV2 [mobilenetv2]. To make these models compatible with evaluated datasets, we follow [ndss26cpmia] and apply several dataset-specific modifications: we adopt ResNet-18 for MNIST and FMNIST, use WideResNet-28 for CIFAR-10 and CIFAR-100, and adjust the input channels of these model architectures to accommodate both grayscale and RGB datasets. We follow the training configurations used in prior work [sp22lira] to mitigate overfitting. Specifically, we use the SGD algorithm with a learning rate of $0.1$, momentum of $0.9$, and weight decay [krogh1991simple] set to $5\times 10^{-4}$. We also adopt a cosine learning rate schedule [loshchilov16sgdr] for optimization and apply data augmentation techniques [aaai20random] during training. Nevertheless, in experiments, we find that some trained models still exhibit overfitting and yield low validation accuracy; in such cases, we rerun these models with different random seeds whenever validation accuracy falls below a threshold. The threshold, along with the accuracy of the target models, is reported in Appendix A.2.
@@
  • LOSS [csf18privacy] uses the loss of the query instance for inference.
  • Entropy [usenix21systematic] leverages a modified prediction entropy estimation for membership inference.
  • Calibration [iclr22calibrate] employs a technique called difficulty calibration, which adjusts the query instance’s loss by calibrating it on shadow models.
  • Attack-R [ccs22enhanced] compares the loss of the query instance on the target model to its losses on shadow models. The final score is the proportion of shadow models for which the target model yields a lower loss.
  • Attack-D [ccs22enhanced] uses model distillation [arxiv15distill] to train shadow models and computes the same ratio scores as Attack-R.
  • SeqMIA [ccs24seqmia] leverages model distillation to replicate the learning process of the target model, extracting various membership metrics from its loss trajectory for attack.
  • LiRA [sp22lira] adapts different strategies for two attack settings. In the adaptive setting, it trains both shadow in and shadow out models and applies a likelihood ratio test for inference. In the non-adaptive setting, it uses only shadow out models to conduct a one-sided hypothesis test.
  • Canary [iclr23canary] improves upon LiRA by employing adversarial optimization to enhance inference performance.
  • GLiRA [tifs25glira] enhances LiRA with model distillation to obtain shadow out models for a one-sided hypothesis test.
  • RMIA [icml24rmia] performs multiple pairwise likelihood ratio tests between the query instance and randomly selected population instances to compute the score.
  • RAPID [ccs24rapid] trains a neural network for membership inference by combining raw and calibrated loss values [iclr22calibrate].
  • PMIA [ndss26cpmia] approximates the posterior odds test by using behaviors of the shadow model’s training set as proxies.
@@
Evaluation Procedures. We divide each dataset into two disjoint subsets: the query set $D_{\text{query}}$ and the auxiliary set $D_{\text{auxiliary}}$. The target model is trained on a randomly sampled subset of $D_{\text{query}}$. In the adaptive setting, we follow [sp22lira, iclr23canary] and provide the adversary with the same data as the query set to prepare their attack (i.e., $D^{\text{adapt}}_{\text{adv}}=D_{\text{query}}$). In the non-adaptive setting, consistent with [ccs24rapid, ndss26cpmia, icml24rmia], the adversary’s dataset is set to the auxiliary set (i.e., $D^{\text{nonadapt}}_{\text{adv}}=D_{\text{auxiliary}}$), ensuring that the adversary cannot access queries when training shadow models. Details about the data split are provided in Table˜18.
Evaluation Metrics. We use the following metrics:
  • TPR@0%FPR. This metric reflects the extent of privacy leakage under the strictest constraint, measuring the highest true positive rate where the attack makes no false positives.
@@
Table 1: Performance comparison of non-adaptive attacks on ResNet across four image datasets (i.e., MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100). $\mathsf{IMIA}$ significantly outperforms the strongest baseline (underlined, PMIA [ndss26cpmia]) while requiring less than 5% of its computational cost (see Section˜4.2 for efficiency comparison). The %Imp. indicates the relative improvement of $\mathsf{IMIA}$ compared to the strongest baseline. The best result is in bold.
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Table 2: Performance comparison of adaptive attacks on ResNet across four image datasets. $\mathsf{IMIA}$ outperforms the strongest baseline (underlined) while requiring less than 5% of its computational cost. The best result is in bold.
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Attack Setup. For baselines, we train the recommended number of shadow models as specified in their papers. For instance, LiRA trains 256 shadow models to estimate the parametric likelihood ratio (see Appendix A.3 for a complete list of shadow model counts). In contrast, $\mathsf{IMIA}$ uses a fixed set of $N=10$ imitative models to highlight its computational efficiency. For imitative training, we first warm up the models for $T_{\text{warmup}}=50$ epochs on MNIST and Fashion-MNIST, and $T_{\text{warmup}}=80$ epochs on CIFAR-10 and CIFAR-100. After the warmup, we train the imitative out models for $T_{1}=100$ epochs across all datasets, matching the number of epochs used in shadow training [sp22lira]. For the imitative in models, we continue training for $T_{2}=20$ using pivot data. For CIFAR-10 and CIFAR-100, we follow [sp22lira, ndss26cpmia, iclr23canary] by querying each instance with $18$ random augmentations to compute the averaged membership score. Regarding hyperparameter configuration, we utilize the default settings from the original baseline implementations without further tuning across all experiments. We acknowledge this as a limitation; while these defaults may not be optimal for every dataset, we maintain this approach due to the prohibitive computational cost required for exhaustive tuning for every baseline. The final results are reported as the average metrics over ten runs.
@@ 4.2 Evaluation of $\mathsf{IMIA}$
Performance in the Non-Adaptive Setting. We first evaluate $\mathsf{IMIA}$ in the non-adaptive setting, and the results on ResNet are shown in Table˜1. $\mathsf{IMIA}$ consistently outperforms all existing methods across the evaluated metrics. For instance, on MNIST, it achieves a true positive rate (TPR) of 1.01% at zero false positive rate (FPR), which is at least three times as high as the best-performing baseline (i.e., PMIA). The advantage is even more pronounced on Fashion-MNIST, with a 2.52% TPR at the same FPR, over ten times higher than baselines. These results represent a substantial advancement, as such datasets were previously considered hard to attack. Although the improvement in balanced accuracy is relatively modest, this is consistent with prior studies [sp22lira, ccs22enhanced, icml24rmia]. Moreover, the consensus in the MIA community is that attacks should be evaluated using TPR at low FPR, where $\mathsf{IMIA}$ shows a clear advantage. Similar trends are observed on other model architectures; results and Receiver Operating Characteristic (ROC) curves are provided in Appendix B.5.
Performance in the Adaptive Setting. We also evaluate $\mathsf{IMIA}$ with SOTA MIAs in the adaptive setting. The results for ResNet are shown in Table˜2, while performance on other architectures can be found in Appenix B.5. $\mathsf{IMIA}$ consistently achieves the best attack performance across all datasets, surpassing strong attacks like LiRA and Canary. For instance, $\mathsf{IMIA}$ achieves nearly double the performance of LiRA on CIFAR-10 at a TPR of 0% FPR, with similar improvements observed across other datasets.
Comparison with Distillation-based MIAs. While the imitative training used in $\mathsf{IMIA}$ is conceptually related to model distillation, it significantly outperforms distillation-based attacks (i.e., Attack-D, SeqMIA, and GLiRA) in both adaptive and non-adaptive settings. For instance, Attack-D yields negligible TPR at low FPRs, and in the non-adaptive setting, $\mathsf{IMIA}$ outperforms SeqMIA and GLiRA by at least a factor of five. We attribute this performance gap to key limitations in existing distillation-based MIAs: (i) These attacks directly apply the standard model distillation process [arxiv15distill], failing to capture the salient membership signals that are important for MIA. (ii) They use distillation to create shadow out models, which can only mimic the target’s behavior on non-member instances. In contrast, $\mathsf{IMIA}$ explicitly caputer membership signals by training both imitative in and out models, leveraging their behavioral discrepancies for a more effective attack.
@@
Efficiency Analysis. We measure the overall runtime of $\mathsf{IMIA}$ in both adaptive and non-adaptive settings on a cluster with a single A100 GPU, comparing it to leading MIAs. We do not include simple baselines (e.g., LOSS attack [csf18privacy]) for comparison, as their performance is far inferior to $\mathsf{IMIA}$. As shown in Table˜3, $\mathsf{IMIA}$ requires substantially less computation. For instance, on CIFAR-10, the SOTA baselines (i.e., PMIA and LiRA) require 6 days to perform their attack, whereas $\mathsf{IMIA}$ completes in less than 7 hours, a reduction of more than 95%. This efficiency is consistently observed across all datasets.
@@
Figure 3: Performance on MNIST under varying computational budgets. $\mathsf{IMIA}$$_{\text{Gaussian}}$ employs a Gaussian likelihood ratio [sp22lira] on imitative models to compute the final scores.
@@
Performance Across Computational Budgets. To evaluate the trade-off between efficiency and attack effectiveness, we benchmark $\mathsf{IMIA}$ against leading MIAs using an A100 GPU in both non-adaptive and adaptive settings. We also introduce $\mathsf{IMIA}$$_{\text{Gaussian}}$, a variant that leverages the same imitative training as $\mathsf{IMIA}$ but employs parametric modeling for inference. Specifically, it follows LiRA [sp22lira] and fits the scaled confidence scores of imitative models as Gaussian distributions, and uses the likelihood ratio as the membership score.
@@ 4.3 Effectiveness of Imitative Training
Does Imitative Training Better Mimic the Target Model? We conduct a set of experiments comparing its ability to mimic the target model’s behavior against that of shadow training. Specifically, we train 256 shadow models and imitative models in the adaptive setting. We then collect the scaled confidence score distributions for both members and non-members by querying the shadow and imitative models. Prior studies [sp22lira, iclr23canary, ndss26cpmia] have empirically validated that these score distributions are well-approximated by a Gaussian distribution. Following this approach, we model each distribution as Gaussian and estimate its mean $\mu$ and standard deviation $\sigma$. Using these parameters, we compute the normalized residual of the target model’s true score $s$ as $r=\frac{s-\mu}{\sigma}$. If the models’ distributions truly replicate the target model’s behavior, the residuals $r$ for both members and non-members should follow a standard normal distribution, i.e., $r\sim N(0,1)$.